# IndicSentEval: How Effectively do Multilingual Transformer Models encode Linguistic Properties for Indic Languages?

### v1

@@ 9 Limitations
The current work focused on only 6 Indic languages. It would be interesting to expand this to more Indic languages. We performed experiments with base versions of various models. It would be interesting to see if larger variants perform better at these tasks and if the trends differ. We experimented with a basic set of linguistic properties. It would be nice to explore more complex linguistic properties like morphological tagging, syntactic parsing, and semantic similarity.
@@ 10 Ethics Statement
All the models used in this work are publicly available on Huggingface and free for research. We utilized publicly accessible resources in SSF format from https://ltrc.iiit.ac.in/showfile.php?filename=downloads/kolhi/ and https://ltrc.iiit.ac.in/showfile.php?filename=downloads/lingResources/newreleases.html. The datasets are licensed under Creative Commons by Non-Commercial 4.0 (CC by NC 4.0). No anticipated risks are associated with using the data from these provided links. We adapted the accessible resources to generate diverse probing and perturbation datasets as required.
Appendix A INDICSENTEVAL dataset statistics
@@
Figure 4: A sample of an SSF formatted sentence in Hindi language.
Appendix B SSF format
Appendix C Details of Multilingual models
@@
Appendix D Probing Results
@@
Appendix E Text Perturbation examples
Tables 14-19 display an illustrative example of text perturbations for each perturbation type per language. For each language, the tables present the original sentence followed by a perturbed sentence resulting from each of the thirteen perturbations. Additionally, English translations are provided for enhanced comprehension.
Table 14: Text perturbation examples for Telugu language.
Table 15: Text perturbation examples for Malayalam language.
Table 16: Text perturbation examples for Hindi language.
Table 17: Text perturbation examples for Marathi language.
Table 18: Text perturbation examples for Kannada language.
Table 19: Text perturbation examples for Urdu language.

### v2

@@ 7 Limitations
The current work focused on only 6 Indic languages. It would be interesting to expand this to more Indic languages.
We performed experiments with base versions of various models. It would be interesting to see if larger variants perform better at these tasks and if the trends differ.
We experimented with a basic set of linguistic properties. It would be nice to explore more complex linguistic properties like morphological tagging, syntactic parsing, and semantic similarity.
@@ 8 Ethics Statement
All the models used in this work are publicly available on Hugging Face and free for research.
@@
We utilized publicly accessible resources in SSF format from https://ltrc.iiit.ac.in/showfile.php?filename=downloads/kolhi/ and https://ltrc.iiit.ac.in/showfile.php?filename=downloads/lingResources/newreleases.html. The datasets are licensed under Creative Commons by Non-Commercial 4.0 (CC BY-NC 4.0). No anticipated risks are associated with using the data from these provided links. We adapted the accessible resources to generate diverse probing and perturbation datasets as required.
Appendix A Related Work
Our work is most closely related to that of Jawahar et al. (2019) and Mohebbi et al. (2021), who focus on understanding the interpretability of language models via probing and observing linguistic structures in English. Our study on multilingual language models for six Indic languages complements their English-focused studies.
Our work also relates to the growing literature that creates resources for low-resource Indic languages. Several recent studies have developed resources including, Indic language models (Khan et al., 2024), Indic NLP suite (Kakwani et al., 2020), Indic BART (Dabre et al., 2021), Indic NLG (Kumar et al., 2022), IndicMT Eval (Dixit et al., 2023), and Naamapadam (Mhaske et al., 2023). Our research is the first to interpret both universal and Indic-specific models for Indic languages through probing representations and assessment of robustness, while also providing implications for future research directions. Overall, we complement these works by studying the linguistic structure of a wide range of low-resource Indic languages on multilingual models.
Non-English and Multilingual Probing Studies: Research on probing language models has increasingly moved beyond English to multilingual and non-English contexts, particularly focusing on morphologically rich and low-resource languages (Pires et al., 2019; Edmiston, 2020; Abdelali et al., 2022; Acs et al., 2024). For instance, Pires et al. (2019) is one of the earliest probing studies on mBERT, showing that mBERT enables zero-shot cross-lingual transfer using shared representations, but performs best on typologically similar languages, struggling with distant or low-resource ones. Extending from previous work, Edmiston (2020) probe BERT’s hidden layers for discrete morphological features in five languages (e.g. gender, tense, case) and observe that BERT encodes discrete morphological features across multiple languages, but lacked explicit analysis of model robustness across language families. More recently, Abdelali et al. (2022) probed several Arabic BERT-based models (and mBERT) and observed that Arabic BERT models capture morphology in lower layers and syntax in higher layers, but had limited focus on dialectal variation and cross-model comparison. Very recently, Acs et al. (2024) introduced a large multilingual probing dataset across 42 languages and found strong morphological encoding in mBERT/XLM-R, but did not analyze model robustness or compare to language-specific models. While these earlier studies primarily focused on morphological encoding in universal multilingual models (like mBERT and XLM-R), IndicSentEval uniquely evaluates both universal multilingual models and Indic-specific models, explicitly comparing their robustness and representational capabilities across six morphologically rich and script-diverse Indic languages.
Critiques of Probing Methodology and Recent Advances Probing methods have faced significant criticism in recent years, primarily for overinterpreting what models “know” from linear classifier performance Hewitt and Liang (2019); Voita and Titov (2020); Ravichander et al. (2021). For instance, Hewitt and Liang (2019) showed that complex probes may "memorize" tasks, not reveal what’s encoded in the representations, ensuring probes extract meaningful, not spurious, information. Extended to previous work, Voita and Titov (2020) proposed using Minimum Description Length (MDL) to evaluate how efficiently linguistic features can be extracted. They show that MDL probes reward simpler models and fewer training samples, improving interpretability and robustness. Further, Ravichander et al. (2021) demonstrate that models can encode linguistic properties even if these properties are not necessary for the task the model was trained on. Specifically, they highlight the importance of careful controls when designing probing experiments, as high probing accuracy does not necessarily indicate that the probed information is utilized by the model during its primary task. In our IndicSentEval work, we address these critiques by: (i) using lightweight probes with standardized architecture across all models and languages; (ii) focusing on relative comparisons (across layers, languages, and perturbations) rather than absolute performance; (iii) including input-level perturbations to test mechanistic interpretability and model robustness; and (iv) complementing standard probing with morpho-semantic tasks tailored to Indic grammar, allowing deeper insight into model behavior for typologically diverse languages.
Differences from English-Centric Findings: Probing studies in non-English and multilingual contexts have revealed notable differences from English-centric findings, due to linguistic diversity and the multilingual training regime Zheng and Liu (2022); Tikhonova et al. (2023); Godunova and Voloshina (2024); Dang et al. (2024). For instance, Zheng and Liu (2022) showed that multilingual models encode universal syntactic features well but struggle with fine-grained morphology (e.g., verb agreement), unlike their performance on English tasks. Similarly, Tikhonova et al. (2023) reported that mBERT lacks the clear layer-wise linguistic hierarchy found in English BERT, suggesting architectural behavior does not generalize across languages. More recently, studies including Godunova and Voloshina (2024) found that discourse-level understanding remains uniformly weak across high- and low-resource languages, contrasting with better performance on lower-level tasks in English. Also, Dang et al. (2024) found that GPT models generalize well to inflect unseen words in morphologically simple languages, but performance declines sharply as morphological complexity increases. In contrast to prior studies that either focused primarily on English or reported high-level trends across many languages, IndicSentEval reveals that while multilingual models exhibit consistent representational hierarchies for English, they show mixed and often unstable behavior across Indic languages, especially in encoding syntactic and morphological features-highlighting language-specific gaps that broader multilingual benchmarks often obscure.
Appendix B SSF format
Figure 4: A sample of an SSF formatted sentence in Hindi language.
Appendix C INDICSENTEVAL dataset statistics
@@
Appendix D Probing Tasks
Surface level tasks (1) Sentence Length (SentLen): Here, the objective is to predict the number of words in sentences, which has been grouped into 8 categories as shown in Table 1 (see main paper).
Appendix E Text Perturbation Examples
Tables 13-18 display an illustrative example of text perturbations for each perturbation type per language. For each language, the tables present the original sentence followed by a perturbed sentence resulting from each of the thirteen perturbations. Additionally, English translations are provided for enhanced comprehension.
Table 13: Text perturbation examples for Telugu language.
Table 14: Text perturbation examples for Malayalam language.
Table 15: Text perturbation examples for Hindi language.
Table 16: Text perturbation examples for Marathi language.
Table 17: Text perturbation examples for Kannada language.
Table 18: Text perturbation examples for Urdu language.
Appendix F Details of Multilingual Models
@@
F.1 Training data proportion of the tested languages
Table 21 shows the proportion of training data for each Indic language within each model’s total Indic language tokens. From Table 21, we make the following observations:
Balanced vs. Imbalanced Distribution:
  • MuRIL model stands out with the most balanced distribution among the six languages shown (12.7% for Hindi, with other languages between 6-9%).
  • In contrast, mGPT shows high Hindi focus (50.5%) with much smaller proportions for other languages.
Hindi Dominance:
  • Most models allocate the largest share of their Indic language tokens to Hindi.
  • This varies dramatically from 12.7% in MuRIL to 50.5% in mGPT.
Coverage Beyond Major Languages:
  • MuRIL allocates 51.1% of its tokens to "Other Indic" languages beyond the six shown, suggesting much broader coverage.
  • This aligns with research showing MuRIL supports 17 Indian languages in total.
Model hi (Hindi) kn (Kannada) ml (Malayalam) mr (Marathi) te (Telugu) ur (Urdu) Other Indic
IndicBERT 24.2% 9.4% 10.1% 7.4% 8.8% 40.1%*
XLM-R 42.9% 4.2% 7.8% 4.4% 6.2% 18.3% 16.2%*
InfoXLM 33.7% 2.8% 9.4% 3.7% 6.5% 8.3% 35.6%*
MuRIL 12.7% 6.4% 7.2% 6.9% 7.0% 8.7% 51.1%*
mT5 41.2% 1.9% 3.1% 24.0% 2.2% 4.1% 23.5%*
mGPT 50.5% 5.0% 5.0% 5.0% 5.0% 9.2% 20.3%*
XGLM 31.4% 4.1% 4.2% 8.5% 6.3% 12.3% 33.2%*
Table 21: Training data proportion of the six Indic languages for both universal and Indic multilingual language models.
Hyper-parameters. We train logistic regression with a regularization parameter CC=20 and use a L2 penalty term. For multi-class tasks, we use the “multinomial” setting while training. All experiments were done on a machine with a T4 GPU.
Appendix G Probing Results
@@
Appendix H Extended Discussion
We evaluated 9 multilingual Transformer-based models on 8 probing tasks to understand linguistic structures in 6 Indic languages, using our contributed IndicSentEval dataset. Indic-specific models like MuRIL and IndicBERT excel due to targeted training, while universal models like mBERT, InfoXLM, BLOOM, mGPT, and XGLM show mixed results. Perturbation analysis reveals decoder-based models are the most robust, as also noted by Neerudu et al. (2023). Verbs and word order are key signals for encoding linguistic structures. TreeDepth is the most sensitive to perturbations, while SubjNum and ObjNum are the most resilient.
Overall, our study represents the first analysis of the interpretability of both Universal and Indic multilingual language models across six Indic languages where several languages have large training corpora while some have less. Our scientific findings from this model interpretability analysis via both probing and perturbations shed light on how language models capture language hierarchy and how training data influences the language understanding across layers in these models. Surprisingly, universal models show greater resilience to perturbations in at least four Indic languages. In contrast, the universal model (mBERT) and the Indic-specific models (IndicBERT and MuRIL) display a more significant accuracy drop across all the Indic languages. This suggests the necessity for multilingual models that are robust to perturbations and can capture language hierarchy regardless of their training data. Overall, our findings demonstrate significant variability in the ability of current multilingual models to capture surface, syntactic, and semantic structures across different Indic languages. A hierarchy of language proficiency is discernible primarily for languages with more extensive training datasets. This highlights the necessity for innovative approaches in multilingual modeling that can accurately capture language structures, even in low-resource languages.